OpenStudy (anonymous):

Need to check my answer: log x + log (x-3) = 1 i got x=1/2(3+-3)

hartnn (hartnn):

how ?

hartnn (hartnn):

doesn't seem correct...

OpenStudy (ghazi):

\[x(x-3)=e^1\]

OpenStudy (anonymous):

law of logs: log a + log b = log (ab) log x + log (x-3) = 1 log x(x-3) = 1 x^2 - 3x = 10 (def. of log says log a = b when 10^b = a) x^2 - 3x - 10 = 0 (x - 5)(x + 2) = 0 x = 5, x = - 2 (not possible b/c domain of log function is x>0) Therefore, x = 5

hartnn (hartnn):

the values should be 5 and -2

OpenStudy (anonymous):

ah great